# Spin-Refined Partition Functions and $\mathcal{CRT}$ Black Holes

### v1

@@ 2.4 Phase diagrams
We are now in a position to construct the phase diagram for $\text{Tr}\big{(}e^{-\beta H+i\tilde{\alpha}J}\big{)}$. The number and nature of the phases depend quite prominently on $\alpha$. We will consider only a few particular physically motivated values of that parameter.
@@
$\text{Tr}\left(e^{-\beta H+i\pi J}\right)=\exp\left(\frac{c\pi^{2}}{12\beta^{2}}\right),$ (40)
@@
which again goes to infinity as $\beta\to 0$. These two parameters $\tilde{\alpha}=\frac{2\pi}{3}$ and $\tilde{\alpha}-\frac{2\pi}{3}$ are mapped into each other by the modular transformation
@@
$F_{\text{rot}}(\beta)\sim-\frac{1}{\beta}.$ (48)
This implies that the partition function with this insertion scales like $e^{\#c}$ at low temperatures.
@@ 2.5 Introducing fermions
$\textrm{Tr}_{\textrm{NS}}\left(e^{-\beta H+i\alpha J}(-1)^{\mathsf{F}}\right)=\textrm{Tr}_{\textrm{NS}}\left(e^{-\beta H+i(\alpha+2\pi)J}\right)$ (49)
@@
$\textrm{Tr}_{\textrm{R}}\left(e^{-\beta H+i\alpha J}\right)=\textrm{Tr}_{\textrm{NS}}\left(e^{\frac{2\pi}{\alpha^{2}+\beta^{2}}\left(-\beta H-i\alpha J\right)}(-1)^{\mathsf{F}}\right).$ (50)
The latter equation states that these two quantities are connected by the S-transform.
@@ 3.5 Even minus odd spin
A similar analysis can be repeated for the microcanonical ensemble. We found 30 distinct solutions $(a(E),r_{+}(E))$, all of which exist at all energies, and numerically computed the entropy $S(E)$ for each one. The results are shown in Figure 8, alongside the entropy of thermal AdS and the $\mathcal{CRT}$-twisted black hole. The latter is just $S_{\mathcal{CRT}}(E)=\frac{1}{2}S_{\text{SAdS}}(E)$, where $S_{\text{SAdS}}(E)$ is given by (60). Of those black hole solutions not excluded by the horizon criterion, one pair of complex saddles dominates the ensemble at all energies. We determined numerically that at high energies, their entropy grows like $\Re(S(E))\sim 2.19E^{3/4}$.
@@ 4 Discussion and further directions
We have investigated the rôle of $\mathcal{CRT}$-twisted black holes in the gravitational path integral and its microstate interpretations. We, in agreement with Benjamin:2024kdg , that they give the leading contribution to the spin-refined partition function $\textrm{Tr}\big{(}e^{-\beta H}R\big{)}$, where $R$ is a reflection, at large temperatures. Thanks to the use of holography, we are able to probe this quantity even at finite temperature. We are also able look more directly into the density of states by working in the microcanonical ensemble. We will review our main results below and point out the questions they raise.
@@ 4.1 The microcanonical ensemble
We have seen that the microcanonical ensemble is not equivalent to the canonical ensemble. At large energies, the former is dominated by complex, rotating black holes whereas at high temperatures the latter is dominated by the $\mathcal{CRT}$-twisted black hole. The reason for the non-equivalence of the ensembles are rapid oscillations in the difference of densities of states for even and odd spins. These oscillations average out to give subdominant contributions to the thermal partition function.
We have checked that this is the case for $\text{AdS}_{3}$ and $\text{AdS}_{5}$. However, it is most likely no true in even-dimensional spacetimes. The reason behind this is essentially just that antipodal map on $S^{D-2}$ is a rotation only for odd $D$. Thus, for even $D$ rotating black holes would not contribute to $\textrm{Tr}_{E}\big{(}e^{-\beta H}R\big{)}$. If we entertain the possibility that there are no other saddles that we are not aware of, then the $\mathcal{CRT}$-twisted black hole (which contributes in any dimension) dominates both the canonical and microcanonical ensembles. This is an important qualitative difference between even and odd dimensions. It would be interesting to see this difference directly in the CFT, for example using the thermal EFT approach of Benjamin:2024kdg .
@@ 4.2 Phase diagrams
Using AdS/CFT, we have been able to probe the phase diagrams for spin-refined partition functions even at finite temperature. This is most clearly achieved in $\text{AdS}_{3}$, where we have a full list of all contributing saddles and thus we can draw the phase diagram for any rotation insertion. Perhaps surprisingly, for rotational insertions we find richer structures than in the case of thermal partition function. In particular, for $\textrm{Tr}\big{(}e^{-\beta H}R\big{)}$ in $\text{AdS}_{3}$ we have three phases. At high temperatures, the dominating saddle is the $\mathcal{CRT}$-twisted black hole. At low temperatures this quantity is dominated by thermal AdS. There is, however, an intermediate regime in which the dominating saddle is a rotating BTZ black hole.
Interestingly, the exchange of the dominance between the latter two occurs at a temperature above the Hawking–Page transition. The points of transition are fixed by modular invariance, and thus we expect this to hold universally, not just for holographic theories. It would be interesting to get more detailed insight into these different phases intrinsically in CFT. Of course, it still remains to be seen whether we can learn more about them on the gravity side. It would be especially interesting to see how the lack of time-orientability of $\mathcal{CRT}$-twisted black holes manifests itself in the $n$-point functions on the boundary. Since that these topological issues are present only behind the horizon, this line of investigation would necessarily lead us outside of the realm of the Euclidean gravity.
We have also found and studied many saddles that potentially contribute to $\textrm{Tr}\big{(}e^{-\beta H}R\big{)}$ in $\text{AdS}_{5}$. The construction of the phase diagram is more complicated here because it is far from obvious which solutions are legitimate saddles that contribute to the gravitational path integral. Based on Figure 7, we expect that no rotating black hole shall ever dominate. [^16] A less desirable alternative would be that the spin-refined density of states is discontinuous in the temperature. On one hand, this is also what happens in $\text{AdS}_{4}$, where the reflection $R$ cannot be reduced to a rotation. On the other hand, we cannot exclude the possibility that there are more exotic saddles that we may have missed. To resolve these ambiguities, it would be beneficial to probe what happens at modest temperatures in the CFT directly. We also hope to find the precise contour-prescription for these quantitites to prove that all of the problematic saddles we have discussed can be rejected from the first principles.
@@ Acknowledgements.
It is a pleasure to thank Daniel Harlow, Don Marolf, Gary Horowitz, Joaquin Turiaci, David Berenstein, and Xiaoyi Liu for helpful discussions. D.G. was supported in part by the Department of Energy under grant DE-SC 0011702. M.K. was supported in part by NSF Grant PHY-2107939.

### v2

@@ 2.4 Phase diagrams
We are now in a position to construct the phase diagram for $\text{Tr}\big{(}e^{-\beta H+i\tilde{\alpha}J}\big{)}$. The number and nature of the phases depend quite prominently on $\alpha$. We will consider only a few particular physically motivated values of that parameter. The phase diagram for arbitrary $\tilde{\alpha}\in\mathbb{Q}$ can be constructed using the phase diagrams discussed in Dijkgraaf:2000fq ; Maloney:2007ud and setting the modular parameter $\tau=\tilde{\alpha}$. In this sense, the results presented here are not new; we are rather reinterpreting them in light of the more recent work of Harlow:2023hjb ; Benjamin:2024kdg .
@@
$\text{Tr}\left(e^{-\beta H+\pi iJ}\right)=\exp\left(\frac{c\pi^{2}}{12\beta^{2}}\right),$ (38)
@@
which again goes to infinity as $\beta\to 0$. The two values of the parameters $\tilde{\alpha}=\frac{2\pi}{3}$ and $\tilde{\alpha}=-\frac{2\pi}{3}$ are mapped into each other by the modular transformation
@@
$F_{\text{rot}}(\beta)\sim-\frac{c}{\beta}.$ (46)
Indeed, Hurwitz’s theorem states that we may approximate any irrational number $\xi$ by a rational one, written in lowest terms as $\frac{m}{n}$, in such a way that
$\left|\xi-\frac{m}{n}\right|<\frac{1}{\sqrt{5}n^{2}}.$ (47)
Now the free energy associated to a sadle $M_{C,D}$ is given by
$F_{\text{rot}}(\tilde{\alpha},\beta)=-\frac{\pi^{2}c}{3C^{2}}\times\frac{1}{\tilde{\beta}^{2}+\left(\xi+\frac{D}{C}\right)^{2}},$ (48)
where $\tilde{\beta}=\frac{\beta}{2\pi}$ and $\xi=\frac{\tilde{\alpha}}{2\pi}$ is any irrational number. By Hurwitz’s theorem, we may choose $C,D$ so that
$\left(\xi+\frac{D}{C}\right)^{2}<\frac{1}{5C^{4}}.$ (49)
It follows that for this choice of $C,D$ we have
$F_{\text{rot}}(\tilde{\alpha},\beta)<-\frac{\pi^{2}c}{3C^{2}}\times\frac{1}{\tilde{\beta}^{2}+\frac{1}{5C^{4}}}.$ (50)
As long as we take $C\sim\tilde{\beta}^{-1/2}$, this quantity is of order $-\frac{c}{\beta}$.
This implies that the partition function with this insertion scales like $e^{\#c}$ at high temperatures. For a similar discussion using continued fractions, see Benjamin:2024kdg .
@@ 2.5 Introducing fermions
$\textrm{Tr}_{\textrm{NS}}\left(e^{-\beta H+i\tilde{\alpha}J}(-1)^{\mathsf{F}}\right)=\textrm{Tr}_{\textrm{NS}}\left(e^{-\beta H+i(\tilde{\alpha}+2\pi)J}\right)$ (51)
@@
$\textrm{Tr}_{\textrm{R}}\left(e^{-\beta H+i\tilde{\alpha}J}\right)=\textrm{Tr}_{\textrm{NS}}\left(e^{\frac{2\pi}{\tilde{\alpha}^{2}+\beta^{2}}\left(-\beta H-i\tilde{\alpha}J\right)}(-1)^{\mathsf{F}}\right).$ (52)
The latter equation states that these two quantities are connected by an S-transformation.
@@ 3.5 Even minus odd spin
A similar analysis can be repeated for the microcanonical ensemble. We found 30 distinct solutions $(a(E),r_{+}(E))$, all of which exist at all energies, and numerically computed the entropy $S(E)$ for each one. The results are shown in Figure 8, alongside the entropy of thermal AdS and the $\mathcal{CRT}$-twisted black hole. The latter is just $S_{\mathcal{CRT}}(E)=\frac{1}{2}S_{\text{SAdS}}(E)$, where $S_{\text{SAdS}}(E)$ is given by (62). Of those black hole solutions not excluded by the horizon criterion, one pair of complex saddles dominates the ensemble at all energies. We determined numerically that at high energies, their entropy grows like $\Re(S(E))\sim 2.19E^{3/4}$. Just as in the 3d case, these dominating saddles also have a nonzero imaginary part of the entropy, leading to a highly oscillatory microcanonical answer for the entropy. Meanwhile, the contribution of $\cal{CRT}$ black holes, while subdominant, is purely real.
@@ 4 Discussion and further directions
We have investigated the rôle of $\mathcal{CRT}$-twisted black holes in the gravitational path integral and its microstate interpretations. In agreement with Benjamin:2024kdg , we find that they give the leading contribution to the spin-refined partition function $\textrm{Tr}\big{(}e^{-\beta H}R\big{)}$, where $R$ is a reflection, at large temperatures. Using gravitational techniques, we are able to probe this quantity even at finite temperature. We are also able look more directly into the density of states by working in the microcanonical ensemble. We will review our main results below and point out some of the questions they raise.
@@ 4.1 The microcanonical ensemble
We have seen that the microcanonical ensemble is not equivalent to the canonical ensemble. [^18] We have, however, used standard thermodynamic relations—which rely on ensemble equivalence—to compute the entropy and action of each saddle if it dominated. This reproduces the correct answers for $\mathcal{CRT}$-twisted and complex rotating black holes, even when they are not dominant. So ensemble equivalence is assumed to hold saddle by saddle, but which saddle dominates depends on the ensemble. At large energies, the former is dominated by complex, rotating black holes, whereas at high temperatures the latter is dominated by the $\mathcal{CRT}$-twisted black hole. The reason for the non-equivalence of the ensembles is the rapid oscillations in the difference of densities of states for even and odd spins. These oscillations average out to give subdominant contributions to the thermal partition function.
We have checked that this is the case for $\text{AdS}_{3}$ and $\text{AdS}_{5}$. However, it is most likely not true in even-dimensional spacetimes. The reason behind this is just that antipodal map on $S^{D-2}$ is a rotation only for odd $D$. Thus, for even $D$ rotating black holes would not contribute to $\textrm{Tr}_{E}\big{(}e^{-\beta H}R\big{)}$. If we entertain the possibility that there are no other saddles that we are not aware of, then the $\mathcal{CRT}$-twisted black hole (which contributes in any dimension) dominates both the canonical and microcanonical ensembles in even dimension. This is an important qualitative difference between even and odd dimensions, and it would be interesting to see it arise directly in the CFT, for example using the thermal EFT approach of Benjamin:2024kdg .
@@ 4.2 Phase diagrams
Using AdS/CFT, it is possible to probe the phase diagrams for spin-refined partition functions even at finite temperature. This is most clearly achieved in $\text{AdS}_{3}$, where we have a full list of all contributing saddles and thus we can draw the phase diagram for any rotational insertion, as was done previously in Dijkgraaf:2000fq ; Maloney:2007ud . Perhaps surprisingly, for such insertions one finds richer structures than in the case of the thermal partition function. For $\textrm{Tr}\big{(}e^{-\beta H}R\big{)}$ in $\text{AdS}_{3}$ we have three phases. At high temperatures, the dominating saddle is the $\mathcal{CRT}$-twisted black hole. At low temperatures this quantity is dominated by thermal AdS. There is, however, an intermediate regime in which the dominating saddle is a rotating BTZ black hole.
Interestingly, the exchange of the dominance between the latter two occurs at a temperature above the Hawking–Page transition. The points of transition are fixed by modular invariance, [^19] We have presented only the simplest example. For the general construction of the phase diagram, see Dijkgraaf:2000fq . and thus we expect this feature to hold universally, not just for holographic theories. It would be interesting to get more detailed insight into these different phases intrinsically in the CFT. Of course, it also remains to be seen whether we can learn more about them on the gravity side. It would be especially interesting to see how the lack of time-orientability of $\mathcal{CRT}$-twisted black holes manifests itself in the $n$-point functions on the boundary. Since these topological issues are present only behind the horizon, this line of investigation would necessarily lead us outside of the realm of the Euclidean gravity.
We have also found and studied many saddles that potentially contribute to $\textrm{Tr}\big{(}e^{-\beta H}R\big{)}$ in $\text{AdS}_{5}$. The construction of the phase diagram is more complicated here because it is far from obvious which solutions are legitimate saddles that contribute to the gravitational path integral. Based on Figure 7, we expect that no rotating black hole shall ever dominate. [^20] A less desirable alternative would be that the spin-refined density of states is discontinuous in $\beta$. On one hand, this is also what happens in $\text{AdS}_{4}$, where the reflection $R$ cannot be reduced to a rotation. On the other hand, we cannot exclude the possibility that there are more exotic saddles that we may have missed. To resolve these ambiguities, it would be beneficial to probe what happens at modest temperatures in the CFT directly. We also hope to find the precise contour-prescription for these quantitites to prove that all of the problematic saddles we have discussed can be rejected from first principles.
@@ Acknowledgements.
It is a pleasure to thank Daniel Harlow, Don Marolf, Gary Horowitz, Joaquin Turiaci, David Berenstein, Xiaoyi Liu, Zixia Wei, Sridip Pal and Paweł Lewulis for helpful discussions. D.G. was supported in part by the Department of Energy under grant DE-SC 0011702. M.K. was supported in part by NSF Grant PHY-2107939.